OpenStudy (azionne):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1) medal

OpenStudy (math&ing001):

Hi, distribute then add like-terms. Distribution: a(b+c)=ab+ac Example: 3(x-2)=3x-2*3=3x-6

OpenStudy (azionne):

can u walk me through the steps its my first time

OpenStudy (math&ing001):

Ok, can you give me distribution for 4(2x-6)

OpenStudy (azionne):

6x-24

OpenStudy (math&ing001):

2*4=8 but I think you got the hang of it.

OpenStudy (azionne):

oops lol

OpenStudy (math&ing001):

Ok, now you can do the same for the other two. Show me what you get =)

TheSmartOne (thesmartone):

So, did you do anymore after where @math&ing001 left you at? @Azionne

OpenStudy (azionne):

No

MsBrains (ms-brains):

x = - 14/11 That is not correct?

OpenStudy (azionne):

No thats right

MsBrains (ms-brains):

Awesome. :)